# Traffic Response Functions: Patterns, Propagation and Congestion

### v1

@@ 4.5 Scaling of responses
The similarities in shape of the responses for smaller time shifts $\tau$ are conspicuous and lead to the question of scalability. Here we try to identify a possible scaling behavior by applying the following approach. Let $\mu_{i}$ be the peak position of the i-th response curve and $\sigma_{i}$ the corresponding full width half maximum. We define
@@
Figure 10 shows examples of velocity responses and their respective rescaling according to Eq. (9) in case of Breitscheid and motorway A3. For the depicted cases the chosen scaling method works well. However, the success of the applied scaling method highly depends on the manifestation of the peaks, or valleys, respectively. In some cases it is not possible to scale our results properly. The determination of the full width half maximum $\sigma_{i}$ is difficult if peaks, or valleys, in the course of responses are not clearly shaped. Nevertheless, the examples depicted in Fig. 10 successfully demonstrate a proof of concept. It is not our goal to carry out a complete analysis of scaling methods and behavior within the scope of this contribution, but the results of our demonstration are satisfying.
@@ 5 Conclusion
Response functions of the fundamental traffic observables velocity, flow and density were analyzed based on congestion indicators for different velocity ranges and spatial subsequent sections on heavily trafficked motorways in NRW, Germany. All three response types of observables clearly showed valleys (hills), or local minima (maxima) within the first 30 minutes of time lags, capturing the impact of congestion and its propagation over several kilometers. We found the correlation between the three fundamental observables reproduced by the responses and showed the importance of considering multiple indicators with different velocity ranges to indicate congestion. We also verified congestion propagating through on- and off-ramps of interchanges into other parts of the motorway network.
@@
Furthermore, via responses, we captured the acceleration-deceleration cycles of traffic participants during congestion, which reflect traffic oscillations, i.e. stop-and-go patterns. Based on this finding we further differentiated congestion patterns with the help of response functions. By considering a sample of the given data we found that the width of the valleys (hills) in the course of the responses coincides with the average period of the occurring traffic oscillations. In addition, we narrowed down congestion phases according to three-phase traffic theory with information gathered from the responses. Motivated by the latter, we derived the propagation velocity of congestion from the responses. Even though the numerical results agree well with empirical observations, it was not possible to derive a generic relation between the propagation velocities and the distance to the indicator section. Lastly, motivated by the similarities in shape of the responses for small time shifts, we successfully conducted an approach for demonstrating scaling behavior of the responses.
@@
The findings of our contribution demonstrate the potential for using response functions as a tool to study traffic dynamics and analyze occurring traffic patterns. Responses reveal the change of traffic observables at given road sections based on prior congestion at other sections. Information of this kind is very useful in applications of traffic engineering, e.g. for estimating, tracking or predicting traffic patterns. A major advantage of the response approach is that it avoids complex calculations and parameter predictions required for the application of various models. Moreover, the presented approach can be conducted with a higher level of details. It is very likely that a more accurate analysis of traffic patterns can be made by determining the responses for each individual lane. The latter is of particular interest with regard to the formation of congestion patterns, their characterization and the determination of their propagation velocity. Possible follow-up studies in this direction seem to be very promising to uncovering further potential of response functions as a useful tool for traffic scientists and engineers.

### v2

@@ 4.5 Scaling of responses
The similarities in shape of the responses for smaller time shifts are striking, prompting the question of scalability. We aim to determine a scaling behavior by applying the following approach. Let $\mu_{i}$ be the peak position of the i-th response curve and $\sigma_{i}$ the corresponding full width at half maximum. We define
@@
Figure 10 provides examples of velocity responses and their respective rescaling in accordance with Eq. (9) for the case of Breitscheid and motorway A3. The selected scaling method proves effective for both cases. Of course the success of the applied scaling method heavely depends on the manifestation of the minima, or maxima, respectively. In some cases it is not possible to properly scale our results. The determination of the full width half maximum $\sigma_{i}$ is difficult if the local extrema are not clearly shaped in the course of the responses. Nevertheless, the examples depicted in Fig. 10 demonstrate a proof of concept and give a first glimpse at universality in the responses.
@@ 5 Conclusion
The response functions of the fundamental traffic observables velocity, flow and density were analyzed based on congestion indicators for different velocity ranges and spatial subsequent sections on heavily trafficked motorways in NRW, Germany. All three response types of observables exhibited clear local minima (maxima) or valleys (hills) within the first 30 minutes of time lags, thereby demonstrating the impact of congestion and its propagation over several kilometers. The correlation between the three fundamental observables is reflected in the responses. Furthermore, the importance of considering multiple indicators with different velocity ranges to indicate congestion was demonstrated. Additionally, the propagation of congestion through on- and off-ramps of interchanges into other parts of the motorway network was verified.
@@
Moreover, the responses capture the acceleration-deceleration cycles during congestion, which reflects traffic oscillations, or stop-and-go patterns. In view of this finding, we proceeded to differentiate between congestion patterns with the aid of response functions. A sample of the data was considered, and it was found that the width of the minima (maxima) or valleys (hills) in the course of the responses coincides with the average period of the occurring traffic oscillations. Furthermore, we narrowed down congestion phases according to three-phase traffic theory, with information gathered from the responses. Additionally, we derived the propagation velocity of congestion from the responses. Despite the numerical results aligning well with empirical observations, it was not feasible to derive a universal relationship between the propagation velocities and the distance to the indicator section. Lastly, motivated by the similarities in the shape of the responses for small time shifts, we successfully conducted an approach for demonstrating a scaling behaviour of the responses.
Our study faced several limitations, primarily related to data quality, detector availability, and spatial constraints. Data quality significantly influenced the observed characteristics and variations, with one-minute intervals in detector measurements limiting our analysis. In addition, differentiating congestion patterns proved challenging; the averaged data limited the ability to analyze the relationship between the widths of minima and traffic oscillations. The determination of propagation velocities and scaling behavior is highly dependent on the distinct manifestations of local minima. These limitations underscore the locally constrained nature of our approach and suggest that improvements in data quality and detector distribution may enhance future analyses.
@@
Despite the limitations, the results of our paper demonstrate that response functions are useful tools to study local traffic dynamics and to analyze emerging traffic patterns. Responses reveal the change in traffic observables at particular road sections based on previous congestion at other sections. Such information is very helpful in traffic engineering applications such as estimating, tracking or predicting traffic patterns. A major advantage of the response approach is that it avoids the complex calculations and parameter predictions required to apply different models. In addition, the presented approach can be carried out with a higher level of detail. It is likely that a more accurate analysis of traffic patterns can be made by determining the responses for each individual lane. This is of particular interest with regard to the formation of congestion patterns, their characterisation and the determination of their propagation velocity. Possible follow-up studies in this direction seem promising to uncover further potential of response functions as a useful tool for traffic scientists and engineers.